Liquid ethanol (C2 H5 OH) is burned with 150% theoretical oxygen in a steady-state, steady-flow process. The reactants enter the combustion chamber at 25 ◦ C and the products leave at 65 ◦ C. The process takes place at 1 bar. Assuming complete combustion (i.e. CO2 and H2 O as products along with any excess O2 ), calculate the heat transfer per kmole of fuel burned.

6.2 Gaseous propane at 25 ◦ C is mixed with air at 400 K and burned. 300% theoretical air is used. What is the adiabatic flame temperature? Again, assume complete combustion.

6.3 Repeat Problems 6.1 and 6.2 using CEA2. Compare the product composition using CEA2 with the assumption of complete combustion. What do you observe?


1
Expert's answer
2021-08-09T08:04:06-0400

"C_2H_5OH + 3O_2 \to 2CO_2 + 3H_2O."


In order to determine the heat transfer from the system a first law

analysis can be performed on the combustion chamber. A general first

law equation for combustion analysis (on a “per mole of fuel basis”) is

"Q_{in} + W_{in} + \Sigma N_r(h^°_f +h + h^°)_r= Q_{out} + W_{out} + \Sigma N_p(h^°_f +h + h^°)_p"


The heat transfer and work will be assumed to be out of the system so

Qin & Win terms are zero. The enthalpies for the reactants will be

determined first.

Note: the circle superscript is used to denote that the

property value is referenced to a 25°C, 1atm reference state.


Reactants

Enthalpy of formation

For stable element O2, the enthalpy of formation is zero.

The enthalpy of formation for C2H5OH(l), is –1235 kJ/mol.


Ideal Gas Enthalpy relative to reference state

In the term, the term is the ideal gas enthalpy at the

temperature of interest and is the ideal gas enthalpy at the reference

temperature. By subtracting from the ideal gas enthalpy at the

temperature of interest is referenced to an ideal gas enthalpy of zero at

the reference state. Doing this for all the reactants and products ensures

that all the enthalpy values are referenced to the same reference state

(25°C, 1atm) and are thus comparable.

Since the reactants are already at the reference state = 0

"h-h^\u00b0 = 0"


Products

Enthalpy of formation

The enthalpy of formations for CO2(g) and H2O(g);

hCO2 = –393520 kJ/kmol

hH2O = –241820 kJ/kmol


Ideal Gas Enthalpy relative to reference state

Since the products are at 65°C, 65°C must first be determined and

then at the reference state of 25°C/298K (this is) must be

subtracted to ensure the enthalpies are all calculated with respect to the

same reference state.


For CO2

at 65°C = 1595 kJ/kmol (Table A-20)

at 25°C= -400 kJ/kmol (Table A-20)

h - h°= 1995 kJ/kmol

For H2O

at 65°C = 1410 kJ/kmol (Table A-23)

at 25°C= -300 kJ/kmol (Table A-23)

h - h° = 1710 kJ/kmol



The specific fuel consumption is not given